# From Tokens to Thoughts: How LLMs and Humans Trade Compression for Meaning

### v4

@@ 6 Discussion and Conclusion
Our information-theoretic investigation reveals a fundamental divergence: LLMs and humans employ starkly different strategies in balancing informational compression with semantic meaning. While LLMs achieve broad categorical alignment with human judgment (RQ1; Section 5.1), they falter in capturing fine-grained semantic nuances such as typicality (RQ2; Section 5.2) and, critically, exhibit vastly different representational efficiency profiles (RQ3; Section 5.3). This pattern strongly suggests that LLMs and humans are optimizing for different objectives.
LLMs appear aggressively optimized for statistical compactness. They form information-theoretic efficient representations, as evidenced by their lower cluster entropy and more “optimal” $\mathcal{L}$ scores. This hints they minimize redundancy and maximize statistical regularity, likely a consequence of their training on immense text corpora. This intense focus on compression, however, limits their capacity to fully encode the rich, prototype-based semantic details vital for deep, human-like understanding.
Human cognition prioritizes adaptive richness, contextual flexibility, and broad functional utility, even if this incurs a cost in statistical compactness as measured by our framework. The higher entropy and $\mathcal{L}$ scores observed for human concepts likely reflect an optimization for a wider array of complex cognitive demands. These include nuanced representations for robust generalization, supporting potent inferential capabilities (causal, functional, goal-oriented), enabling effective communication through learnable and shareable structures, and grounding concepts in rich, multimodal experiences. The brain’s neural architecture itself may inherently favor distributed, context-sensitive, and adaptable representations over statically optimal compression. Human cognition, therefore, appears to “invest” in what our statistical measures register as inefficiency for better adaptability and versatility.
The noteworthy performance of smaller encoder models like BERT in specific alignment tasks (Section 5.1) also underscores that architectural design and pre-training objectives significantly influence a model’s ability to abstract human-like conceptual information. This observation highlights important avenues for future AI development focused on enhancing human-AI alignment.
These divergent representational strategies carry significant implications. For AI development, achieving more human-like understanding demands moving beyond current paradigms often centered on scaling and statistical pattern matching. Future efforts should explore principles that explicitly foster richer, more nuanced conceptual structures; our information-theoretic framework and $\mathcal{L}$ objective (Section 4) offer a potential class of tools for guiding and evaluating models toward this more human-like balance. For cognitive science, LLMs, with their distinct optimization biases, serve as valuable computational foils. Comparing their operational strategies against human performance can illuminate the unique constraints and multifaceted objectives that have shaped human concept formation, providing a powerful testbed for cognitive theories.
In essence, LLMs excel at statistical compressibility, treading a representational path fundamentally distinct from human cognition, which champions adaptive richness and functional utility, often above sheer statistical efficiency. This core difference is critical: it highlights current limitations in AI’s pursuit of human-like understanding and charts vital directions for future research. Progressing AI “from tokens to thoughts“, towards systems that genuinely comprehend and reason, will necessitate embracing principles that cultivate this richer, contextually-aware conceptual structure. Our framework offers a quantitative step in this direction, encouraging further exploration of how apparent “inefficiencies” might, in fact, be hallmarks of robust, human-like intelligence.
Appendix A Limitations
@@
A.1 Dataset Access Details
The aggregated and digitized human categorization datasets from Rosch [1973a, 1975], McCloskey and Glucksberg [1978] are made available in CSV format at: https://huggingface.co/datasets/CShani/human-concepts.
A.2 LLM Details
  • BERT family: deberta-large, bert-large-uncased, roberta-large [Devlin et al., 2019, He et al., 2020, Zhuang et al., 2021].
  • QWEN family: qwen2-0.5b, qwen2.5-0.5b, qwen1.5-0.5b, qwen2.5-1.5b, qwen2-1.5b, qwen1.5-1.5b, qwen1.5-4b, qwen2.5-4b, qwen2-7b, qwen1.5-14b, qwen1.5-32b, qwen1.5-72b, qwen2.5-72b [Bai et al., 2023, Yang et al., 2024].
  • Llama family: llama-3.2-1b, llama-3.1-8b, llama-3-8b, llama-3-70b, llama-3.1-70b [Touvron et al., 2023a, b, Grattafiori et al., 2024].
  • Phi family: phi-1.5, phi-1, phi-2, phi-4 [Javaheripi et al., 2023, Abdin et al., 2024, Abouelenin et al., 2025].
  • Gemma family: gemma-2b, gemma-2-2b, gemma-7b, gemma-2-9b [Team et al., 2024, 2025].
  • Mistral family: mistral-7b-v0.3 [Karamcheti et al., 2021].
  • GPT family: gpt2, gpt2-medium [Radford et al., 2019].
  • OLMo family: olmo-7b [Groeneveld et al., 2024].
  • ViT family: Clip ViT-B/32, Clip ViT-B/16 [Radford et al., 2021].
  • Classic static embeddings: GloVe, Word2Vec [Pennington et al., 2014, Mikolov et al., 2013a, b].
A.3 Contextual Prompts and Pooling Strategies
@@
Figure 5: The neutral prompt template "This is a word. " demonstrates balanced performance and moderate consistency, making it ideal for baseline comparisons. Performance Distribution across various prompts.
@@
A.4 Static vs. Contextual AMI Exploration
@@ A.7 Tokenizer Analysis
  • Qwen: Qwen1.5-0.5B (representative), Qwen1.5-1.8B, Qwen1.5-14B, Qwen2-0.5B, Qwen2-7B, Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-32B, Qwen1.5-32B, phi-4111phi-4 has a different tokenizer than the rest of Phi family. The results of its tokenizer match the tokenizer of the Qwen family.
A.8 Additional Clustering Metrics
@@
A.9 Detailed AMI Scores per Model and Dataset
@@
[Rosch, 1973c] bert-large-uncased 0.19453 0.2011 0.11336
[Rosch, 1975] bert-large-uncased 0.16547 0.27324 0.2216
[McCloskey and Glucksberg, 1978] bert-large-uncased 0.12003 0.15934 0.06306
[Rosch, 1973c] FacebookAI/roberta-large 0.1021 0.10666 0.03393
[Rosch, 1975] FacebookAI/roberta-large 0.12138 0.23938 0.14165
[McCloskey and Glucksberg, 1978] FacebookAI/roberta-large 0.06271 0.08873 0.03173
[Rosch, 1973c] google-t5/t5-large 0.16583 0.16855 0.03676
[Rosch, 1975] google-t5/t5-large -0.03799 0.04179 0.00758
[McCloskey and Glucksberg, 1978] google-t5/t5-large 0.06146 0.08825 0.0082
[Rosch, 1973c] google/gemma-2-27b 0.08523 0.09065 0.04158
[Rosch, 1975] google/gemma-2-27b 0.04276 0.10062 0.06244
[McCloskey and Glucksberg, 1978] google/gemma-2-27b 0.07814 0.10274 0.04364
[Rosch, 1973c] google/gemma-2-2b 0.04029 0.04107 0.01212
[Rosch, 1975] google/gemma-2-2b 0.04529 0.14844 0.07596
[McCloskey and Glucksberg, 1978] google/gemma-2-2b 0.09953 0.13593 0.06326
[Rosch, 1973c] google/gemma-2-9b 0.1222 0.12757 0.06053
[Rosch, 1975] google/gemma-2-9b 0.07841 0.16126 0.09617
[McCloskey and Glucksberg, 1978] google/gemma-2-9b 0.10879 0.13997 0.06439
[Rosch, 1973c] google/gemma-2b 0.04336 0.04616 0.01593
[Rosch, 1975] google/gemma-2b -0.00353 0.04483 0.01577
[McCloskey and Glucksberg, 1978] google/gemma-2b 0.03472 0.05484 0.02142
[Rosch, 1973c] google/gemma-7b 0.04459 0.04547 0.01052
[Rosch, 1975] google/gemma-7b -0.03055 0.02644 0.01506
[McCloskey and Glucksberg, 1978] google/gemma-7b 0.03338 0.05724 0.02176
[Rosch, 1973c] meta-llama/Llama-3.1-70B 0.03008 0.03528 0.01936
[Rosch, 1975] meta-llama/Llama-3.1-70B -0.07026 0.02636 0.00392
[McCloskey and Glucksberg, 1978] meta-llama/Llama-3.1-70B -0.04773 0.00972 0.00236
[Rosch, 1973c] meta-llama/Llama-3.1-8B 0.00473 0.00393 0.00023
[Rosch, 1975] meta-llama/Llama-3.1-8B -0.03928 0.05489 0.01884
[McCloskey and Glucksberg, 1978] meta-llama/Llama-3.1-8B -0.02671 0.02208 6.00E-05
[Rosch, 1973c] meta-llama/Llama-3.2-1B 0.01936 0.01567 0.00246
[Rosch, 1975] meta-llama/Llama-3.2-1B -0.01876 0.05663 0.00782
[McCloskey and Glucksberg, 1978] meta-llama/Llama-3.2-1B 0.03625 0.06798 0.01352
[Rosch, 1973c] meta-llama/Llama-3.2-3B 0.03757 0.03537 0.00876
[Rosch, 1975] meta-llama/Llama-3.2-3B 0.01893 0.09619 0.03193
[McCloskey and Glucksberg, 1978] meta-llama/Llama-3.2-3B 0.03914 0.07395 0.0202
[Rosch, 1973c] meta-llama/Meta-Llama-3-70B 0.02289 0.03133 0.01514
[Rosch, 1975] meta-llama/Meta-Llama-3-70B -0.06428 0.0185 0.00554
[McCloskey and Glucksberg, 1978] meta-llama/Meta-Llama-3-70B -0.04595 0.01068 0.00272
[Rosch, 1973c] meta-llama/Meta-Llama-3-8B 0.03512 0.02852 0.00225
[Rosch, 1975] meta-llama/Meta-Llama-3-8B -0.06011 0.03694 0.00676
[McCloskey and Glucksberg, 1978] meta-llama/Meta-Llama-3-8B -0.0355 0.0219 0.00676
[Rosch, 1973c] microsoft/deberta-large 0.03748 0.03909 0.01467
[Rosch, 1975] microsoft/deberta-large 0.16568 0.28993 0.20527
[McCloskey and Glucksberg, 1978] microsoft/deberta-large 0.03217 0.06175 0.03019
[Rosch, 1973c] microsoft/phi-1_5 0.02102 0.01786 0.0075
[Rosch, 1975] microsoft/phi-1_5 0.03989 0.13887 0.04305
[McCloskey and Glucksberg, 1978] microsoft/phi-1_5 0.00895 0.05215 0.00639
[Rosch, 1973c] microsoft/phi-1 0.0249 0.01698 0.00133
[Rosch, 1975] microsoft/phi-1 -0.03625 0.02811 0.00217
[McCloskey and Glucksberg, 1978] microsoft/phi-1 -0.01148 0.03085 0.00371
[Rosch, 1973c] microsoft/phi-2 0.03703 0.02968 0.00404
[Rosch, 1975] microsoft/phi-2 -0.03654 0.04227 0.03942
[McCloskey and Glucksberg, 1978] microsoft/phi-2 -0.00254 0.02531 0.00533
[Rosch, 1973c] microsoft/phi-4 0.03075 0.03043 0.01076
[Rosch, 1975] microsoft/phi-4 -0.06737 0.00092 -0.01361
[McCloskey and Glucksberg, 1978] microsoft/phi-4 -0.01789 0.02705 0.00066
[Rosch, 1973c] mistralai/Mistral-7B-v0.3 0.0425 0.03507 0.00357
[Rosch, 1975] mistralai/Mistral-7B-v0.3 -0.05018 0.01217 0.0177
[McCloskey and Glucksberg, 1978] mistralai/Mistral-7B-v0.3 -0.01264 0.03902 0.00931
[Rosch, 1973c] Qwen/Qwen1.5-0.5B 0.00148 -0.00225 0.00399
[Rosch, 1975] Qwen/Qwen1.5-0.5B -0.01538 0.04833 0.0095
[McCloskey and Glucksberg, 1978] Qwen/Qwen1.5-0.5B 0.02559 0.06023 0.00771
[Rosch, 1973c] Qwen/Qwen1.5-1.8B 0.03397 0.03232 0.01034
[Rosch, 1975] Qwen/Qwen1.5-1.8B -0.01129 0.05803 0.00683
[McCloskey and Glucksberg, 1978] Qwen/Qwen1.5-1.8B -0.00541 0.03614 0.00538
[Rosch, 1973c] Qwen/Qwen1.5-14B 0.0372 0.02738 0.0028
[Rosch, 1975] Qwen/Qwen1.5-14B -0.02604 0.05153 0.01211
[McCloskey and Glucksberg, 1978] Qwen/Qwen1.5-14B 0.00124 0.04136 0.00338
[Rosch, 1973c] Qwen/Qwen1.5-32B 0.02638 0.02436 0.00409
[Rosch, 1975] Qwen/Qwen1.5-32B -0.03413 0.02526 -0.00665
[McCloskey and Glucksberg, 1978] Qwen/Qwen1.5-32B -0.01991 0.02124 -0.00059
[Rosch, 1973c] Qwen/Qwen1.5-4B 0.03803 0.04058 0.01742
[Rosch, 1975] Qwen/Qwen1.5-4B -0.03309 0.03988 0.01678
[McCloskey and Glucksberg, 1978] Qwen/Qwen1.5-4B -0.03997 0.00548 -0.00028
[Rosch, 1973c] Qwen/Qwen1.5-72B 0.03697 0.02892 0.00144
[Rosch, 1975] Qwen/Qwen1.5-72B -0.06184 0.02213 0.0017
[McCloskey and Glucksberg, 1978] Qwen/Qwen1.5-72B -0.02022 0.02918 0.00297
[Rosch, 1973c] Qwen/Qwen2-0.5B 0.02266 0.01923 0.00662
[Rosch, 1975] Qwen/Qwen2-0.5B 0.0515 0.14571 0.04999
[McCloskey and Glucksberg, 1978] Qwen/Qwen2-0.5B 0.01508 0.04357 0.00643
[Rosch, 1973c] Qwen/Qwen2-1.5B 0.02956 0.02779 0.00544
[Rosch, 1975] Qwen/Qwen2-1.5B -0.03595 0.03443 -0.01099
[McCloskey and Glucksberg, 1978] Qwen/Qwen2-1.5B 0.01768 0.05407 0.01604
[Rosch, 1973c] Qwen/Qwen2-7B 0.06424 0.06439 0.02067
[Rosch, 1975] Qwen/Qwen2-7B 0.0333 0.09155 0.02832
[McCloskey and Glucksberg, 1978] Qwen/Qwen2-7B 0.05329 0.07599 0.01977
[Rosch, 1973c] Qwen/Qwen2.5-0.5B 0.03165 0.03291 0.01029
[Rosch, 1975] Qwen/Qwen2.5-0.5B -0.06534 -0.0196 -0.01165
[McCloskey and Glucksberg, 1978] Qwen/Qwen2.5-0.5B 0.0062 0.04191 0.0054
[Rosch, 1973c] Qwen/Qwen2.5-1.5B 0.04838 0.0489 0.0129
[Rosch, 1975] Qwen/Qwen2.5-1.5B 0.03785 0.113 0.02761
[McCloskey and Glucksberg, 1978] Qwen/Qwen2.5-1.5B 0.06166 0.08675 0.03162
[Rosch, 1973c] Qwen/Qwen2.5-3B 0.03882 0.0348 0.00465
[Rosch, 1975] Qwen/Qwen2.5-3B 0.03977 0.10821 0.04302
[McCloskey and Glucksberg, 1978] Qwen/Qwen2.5-3B 0.03416 0.07307 0.02959
[Rosch, 1973c] Qwen/Qwen2.5-7B 0.0529 0.05051 0.01605
[Rosch, 1975] Qwen/Qwen2.5-7B -0.00905 0.03227 0.01044
[McCloskey and Glucksberg, 1978] Qwen/Qwen2.5-7B 0.00222 0.02759 0.00551
A.10 Correlation between Human Typicality Judgments and LLM Internal Cluster Geometry

### v5

@@ 6 Discussion and Conclusion
We investigated how LLMs and humans navigate the compression-meaning trade-off in conceptual representation. Using information-theoretic analysis of 40+ models against classic cognitive benchmarks, we reveal fundamental differences in their representational strategies.
Key Findings. LLMs achieve broad categorical alignment with humans (AMI $\approx$ 0.55), successfully partitioning semantic space into recognizable categories. However, they fail to capture the internal structure that makes these categories cognitively useful, as typicality correlations remain weak ($\rho<0.2$) across model families. Most strikingly, when evaluated on the compression-meaning trade-off, LLMs consistently achieve lower (better) $\mathcal{L}$ scores than human categories, indicating they optimize for statistical efficiency over semantic richness. This pattern holds across architectures, though encoder models surprisingly outperform decoder models in human alignment despite being orders of magnitude smaller. Training dynamics analysis reveals rapid category formation followed by architectural reorganization that shifts semantic processing from deep to mid-network layers, suggesting efficiency optimization continues throughout training.
Implications. These findings challenge the assumption that statistical optimality equals understanding. LLMs excel at their training objective, which is minimizing prediction error, but this drives them toward representations that sacrifice semantic nuances. Encoder models’ superior alignment with human representations questions the current paradigm of unified, scaled decoder models, suggesting that language understanding and generation may require distinct architectures and rely on different processes. Our framework provides quantitative tools for monitoring the compression-meaning balance in future systems.
Conclusions. Our findings reveal an apparent paradox, showing that LLMs are simultaneously better and worse than humans. This occurs because LLMs and humans employ divergent strategies: statistical compression versus semantic richness, likely reflecting different optimization pressures. While LLMs process billions of tokens efficiently, humans enable flexible reasoning and generalization. Progress toward human-like AI may require preserving the apparent “inefficiencies” that support cognitive flexibility. We provide theoretical understanding, practical metrics, and high-quality digital benchmarks to develop more human-aligned representations. We encourage the community to utilize the data and metrics for future research towards making AI more human-like.
7 Ethics statement
Our study relies exclusively on publicly available LLMs and digitized datasets from classic cognitive psychology experiments (Rosch, 1973a; 1975; McCloskey & Glucksberg, 1978). No new human subject data was collected, and all benchmark data we release have been properly attributed and curated to preserve research integrity.
We do not foresee privacy, security, or fairness risks arising from our analyses. Our contribution is methodological and theoretical, focusing on representational trade-offs between humans and LLMs. Nevertheless, we acknowledge that insights into model-human divergences could influence how future systems are designed. We caution that optimizing solely for statistical efficiency without considering semantic richness may exacerbate risks of misinterpretation or oversimplification in socially sensitive applications.
We declare no conflicts of interest or external sponsorship that could bias the reported findings.
8 Reproducibility statement
We have taken several steps to ensure reproducibility. All digitized human categorization datasets used in our analyses are publicly released in machine-readable form (Appendix B.1). Detailed model specifications, including architectures, scales, and hyperparameters, are provided in Appendix B.2, and we document embedding extraction procedures, pooling strategies, and prompt templates in Appendix B.3-B.4. Full experimental results, including layer-wise analyses, clustering metrics, and training dynamics across checkpoints, are reported in the appendices (B.5-B.14). Our theoretical framework and derivations are described in Section 4, with complete definitions and formulations provided to enable replication. We will release the code for dataset processing, embedding extraction, and evaluation upon acceptance (to preserve anonymity).
Appendix A Appendix
Appendix B Limitations
@@
B.1 Dataset Access Details
The aggregated and digitized human categorization datasets from Rosch (1973a; 1975); McCloskey & Glucksberg (1978) are made available in CSV format at: [URL deduced for anonymity; Data is attached as Supplementary Material].
B.2 LLM Details
  • BERT family: deberta-large, bert-large-uncased, roberta-large (Devlin et al., 2019; He et al., 2020; Zhuang et al., 2021).
  • QWEN family: qwen2-0.5b, qwen2.5-0.5b, qwen1.5-0.5b, qwen2.5-1.5b, qwen2-1.5b, qwen1.5-1.5b, qwen1.5-4b, qwen2.5-4b, qwen2-7b, qwen1.5-14b, qwen1.5-32b, qwen1.5-72b, qwen2.5-72b (Bai et al., 2023; Yang et al., 2024).
  • Llama family: llama-3.2-1b, llama-3.1-8b, llama-3-8b, llama-3-70b, llama-3.1-70b (Touvron et al., 2023a; b; Grattafiori et al., 2024).
  • Phi family: phi-1.5, phi-1, phi-2, phi-4 (Javaheripi et al., 2023; Abdin et al., 2024; Abouelenin et al., 2025).
  • Gemma family: gemma-2b, gemma-2-2b, gemma-7b, gemma-2-9b (Team et al., 2024; 2025).
  • Mistral family: mistral-7b-v0.3 (Karamcheti et al., 2021).
  • GPT family: gpt2, gpt2-medium (Radford et al., 2019).
  • OLMo family: olmo-7b (Groeneveld et al., 2024).
  • ViT family: Clip ViT-B/32, Clip ViT-B/16 (Radford et al., 2021).
  • Classic static embeddings: GloVe, Word2Vec (Pennington et al., 2014; Mikolov et al., 2013a; b).
B.3 Contextual Prompts and Pooling Strategies
@@
Figure 5: The neutral prompt template ”This is a word. ” demonstrates balanced performance and moderate consistency, making it ideal for baseline comparisons. Performance Distribution across various prompts.
@@
B.4 Static vs. Contextual AMI Exploration
@@ B.5 Training Dynamics
Figure 8: OLMo-7B develops conceptual structure through two-phase dynamics. Left: AMI with human categories rises rapidly in early training then refines gradually. Right: Peak semantic processing migrates from deep (layer 28) toward mid-network layers during training, revealing architectural reorganization for efficiency. Representative checkpoints shown; full 57-checkpoint analysis in Appendix B.5.
@@ B.7 Tokenizer Analysis
  • Qwen: Qwen1.5-0.5B (representative), Qwen1.5-1.8B, Qwen1.5-14B, Qwen2-0.5B, Qwen2-7B, Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-32B, Qwen1.5-32B, phi-4222phi-4 has a different tokenizer than the rest of Phi family. The results of its tokenizer match the tokenizer of the Qwen family.
B.8 Additional Clustering Metrics
@@
B.9 Detailed AMI Scores per Model and Dataset
@@
(Rosch, 1973c) bert-large-uncased 0.19453 0.2011 0.11336
(Rosch, 1975) bert-large-uncased 0.16547 0.27324 0.2216
(McCloskey & Glucksberg, 1978) bert-large-uncased 0.12003 0.15934 0.06306
(Rosch, 1973c) FacebookAI/roberta-large 0.1021 0.10666 0.03393
(Rosch, 1975) FacebookAI/roberta-large 0.12138 0.23938 0.14165
(McCloskey & Glucksberg, 1978) FacebookAI/roberta-large 0.06271 0.08873 0.03173
(Rosch, 1973c) google-t5/t5-large 0.16583 0.16855 0.03676
(Rosch, 1975) google-t5/t5-large -0.03799 0.04179 0.00758
(McCloskey & Glucksberg, 1978) google-t5/t5-large 0.06146 0.08825 0.0082
(Rosch, 1973c) google/gemma-2-27b 0.08523 0.09065 0.04158
(Rosch, 1975) google/gemma-2-27b 0.04276 0.10062 0.06244
(McCloskey & Glucksberg, 1978) google/gemma-2-27b 0.07814 0.10274 0.04364
(Rosch, 1973c) google/gemma-2-2b 0.04029 0.04107 0.01212
(Rosch, 1975) google/gemma-2-2b 0.04529 0.14844 0.07596
(McCloskey & Glucksberg, 1978) google/gemma-2-2b 0.09953 0.13593 0.06326
(Rosch, 1973c) google/gemma-2-9b 0.1222 0.12757 0.06053
(Rosch, 1975) google/gemma-2-9b 0.07841 0.16126 0.09617
(McCloskey & Glucksberg, 1978) google/gemma-2-9b 0.10879 0.13997 0.06439
(Rosch, 1973c) google/gemma-2b 0.04336 0.04616 0.01593
(Rosch, 1975) google/gemma-2b -0.00353 0.04483 0.01577
(McCloskey & Glucksberg, 1978) google/gemma-2b 0.03472 0.05484 0.02142
(Rosch, 1973c) google/gemma-7b 0.04459 0.04547 0.01052
(Rosch, 1975) google/gemma-7b -0.03055 0.02644 0.01506
(McCloskey & Glucksberg, 1978) google/gemma-7b 0.03338 0.05724 0.02176
(Rosch, 1973c) meta-llama/Llama-3.1-70B 0.03008 0.03528 0.01936
(Rosch, 1975) meta-llama/Llama-3.1-70B -0.07026 0.02636 0.00392
(McCloskey & Glucksberg, 1978) meta-llama/Llama-3.1-70B -0.04773 0.00972 0.00236
(Rosch, 1973c) meta-llama/Llama-3.1-8B 0.00473 0.00393 0.00023
(Rosch, 1975) meta-llama/Llama-3.1-8B -0.03928 0.05489 0.01884
(McCloskey & Glucksberg, 1978) meta-llama/Llama-3.1-8B -0.02671 0.02208 6.00E-05
(Rosch, 1973c) meta-llama/Llama-3.2-1B 0.01936 0.01567 0.00246
(Rosch, 1975) meta-llama/Llama-3.2-1B -0.01876 0.05663 0.00782
(McCloskey & Glucksberg, 1978) meta-llama/Llama-3.2-1B 0.03625 0.06798 0.01352
(Rosch, 1973c) meta-llama/Llama-3.2-3B 0.03757 0.03537 0.00876
(Rosch, 1975) meta-llama/Llama-3.2-3B 0.01893 0.09619 0.03193
(McCloskey & Glucksberg, 1978) meta-llama/Llama-3.2-3B 0.03914 0.07395 0.0202
(Rosch, 1973c) meta-llama/Meta-Llama-3-70B 0.02289 0.03133 0.01514
(Rosch, 1975) meta-llama/Meta-Llama-3-70B -0.06428 0.0185 0.00554
(McCloskey & Glucksberg, 1978) meta-llama/Meta-Llama-3-70B -0.04595 0.01068 0.00272
(Rosch, 1973c) meta-llama/Meta-Llama-3-8B 0.03512 0.02852 0.00225
(Rosch, 1975) meta-llama/Meta-Llama-3-8B -0.06011 0.03694 0.00676
(McCloskey & Glucksberg, 1978) meta-llama/Meta-Llama-3-8B -0.0355 0.0219 0.00676
(Rosch, 1973c) microsoft/deberta-large 0.03748 0.03909 0.01467
(Rosch, 1975) microsoft/deberta-large 0.16568 0.28993 0.20527
(McCloskey & Glucksberg, 1978) microsoft/deberta-large 0.03217 0.06175 0.03019
(Rosch, 1973c) microsoft/phi-1_5 0.02102 0.01786 0.0075
(Rosch, 1975) microsoft/phi-1_5 0.03989 0.13887 0.04305
(McCloskey & Glucksberg, 1978) microsoft/phi-1_5 0.00895 0.05215 0.00639
(Rosch, 1973c) microsoft/phi-1 0.0249 0.01698 0.00133
(Rosch, 1975) microsoft/phi-1 -0.03625 0.02811 0.00217
(McCloskey & Glucksberg, 1978) microsoft/phi-1 -0.01148 0.03085 0.00371
(Rosch, 1973c) microsoft/phi-2 0.03703 0.02968 0.00404
(Rosch, 1975) microsoft/phi-2 -0.03654 0.04227 0.03942
(McCloskey & Glucksberg, 1978) microsoft/phi-2 -0.00254 0.02531 0.00533
(Rosch, 1973c) microsoft/phi-4 0.03075 0.03043 0.01076
(Rosch, 1975) microsoft/phi-4 -0.06737 0.00092 -0.01361
(McCloskey & Glucksberg, 1978) microsoft/phi-4 -0.01789 0.02705 0.00066
(Rosch, 1973c) mistralai/Mistral-7B-v0.3 0.0425 0.03507 0.00357
(Rosch, 1975) mistralai/Mistral-7B-v0.3 -0.05018 0.01217 0.0177
(McCloskey & Glucksberg, 1978) mistralai/Mistral-7B-v0.3 -0.01264 0.03902 0.00931
(Rosch, 1973c) Qwen/Qwen1.5-0.5B 0.00148 -0.00225 0.00399
(Rosch, 1975) Qwen/Qwen1.5-0.5B -0.01538 0.04833 0.0095
(McCloskey & Glucksberg, 1978) Qwen/Qwen1.5-0.5B 0.02559 0.06023 0.00771
(Rosch, 1973c) Qwen/Qwen1.5-1.8B 0.03397 0.03232 0.01034
(Rosch, 1975) Qwen/Qwen1.5-1.8B -0.01129 0.05803 0.00683
(McCloskey & Glucksberg, 1978) Qwen/Qwen1.5-1.8B -0.00541 0.03614 0.00538
(Rosch, 1973c) Qwen/Qwen1.5-14B 0.0372 0.02738 0.0028
(Rosch, 1975) Qwen/Qwen1.5-14B -0.02604 0.05153 0.01211
(McCloskey & Glucksberg, 1978) Qwen/Qwen1.5-14B 0.00124 0.04136 0.00338
(Rosch, 1973c) Qwen/Qwen1.5-32B 0.02638 0.02436 0.00409
(Rosch, 1975) Qwen/Qwen1.5-32B -0.03413 0.02526 -0.00665
(McCloskey & Glucksberg, 1978) Qwen/Qwen1.5-32B -0.01991 0.02124 -0.00059
(Rosch, 1973c) Qwen/Qwen1.5-4B 0.03803 0.04058 0.01742
(Rosch, 1975) Qwen/Qwen1.5-4B -0.03309 0.03988 0.01678
(McCloskey & Glucksberg, 1978) Qwen/Qwen1.5-4B -0.03997 0.00548 -0.00028
(Rosch, 1973c) Qwen/Qwen1.5-72B 0.03697 0.02892 0.00144
(Rosch, 1975) Qwen/Qwen1.5-72B -0.06184 0.02213 0.0017
(McCloskey & Glucksberg, 1978) Qwen/Qwen1.5-72B -0.02022 0.02918 0.00297
(Rosch, 1973c) Qwen/Qwen2-0.5B 0.02266 0.01923 0.00662
(Rosch, 1975) Qwen/Qwen2-0.5B 0.0515 0.14571 0.04999
(McCloskey & Glucksberg, 1978) Qwen/Qwen2-0.5B 0.01508 0.04357 0.00643
(Rosch, 1973c) Qwen/Qwen2-1.5B 0.02956 0.02779 0.00544
(Rosch, 1975) Qwen/Qwen2-1.5B -0.03595 0.03443 -0.01099
(McCloskey & Glucksberg, 1978) Qwen/Qwen2-1.5B 0.01768 0.05407 0.01604
(Rosch, 1973c) Qwen/Qwen2-7B 0.06424 0.06439 0.02067
(Rosch, 1975) Qwen/Qwen2-7B 0.0333 0.09155 0.02832
(McCloskey & Glucksberg, 1978) Qwen/Qwen2-7B 0.05329 0.07599 0.01977
(Rosch, 1973c) Qwen/Qwen2.5-0.5B 0.03165 0.03291 0.01029
(Rosch, 1975) Qwen/Qwen2.5-0.5B -0.06534 -0.0196 -0.01165
(McCloskey & Glucksberg, 1978) Qwen/Qwen2.5-0.5B 0.0062 0.04191 0.0054
(Rosch, 1973c) Qwen/Qwen2.5-1.5B 0.04838 0.0489 0.0129
(Rosch, 1975) Qwen/Qwen2.5-1.5B 0.03785 0.113 0.02761
(McCloskey & Glucksberg, 1978) Qwen/Qwen2.5-1.5B 0.06166 0.08675 0.03162
(Rosch, 1973c) Qwen/Qwen2.5-3B 0.03882 0.0348 0.00465
(Rosch, 1975) Qwen/Qwen2.5-3B 0.03977 0.10821 0.04302
(McCloskey & Glucksberg, 1978) Qwen/Qwen2.5-3B 0.03416 0.07307 0.02959
(Rosch, 1973c) Qwen/Qwen2.5-7B 0.0529 0.05051 0.01605
(Rosch, 1975) Qwen/Qwen2.5-7B -0.00905 0.03227 0.01044
(McCloskey & Glucksberg, 1978) Qwen/Qwen2.5-7B 0.00222 0.02759 0.00551
B.10 Correlation between Human Typicality Judgments and LLM Internal Cluster Geometry